# Learning Mean-Field Games through Mean-Field Actor-Critic Flow

### v1

@@ Appendix C Proofs for the critic
Substituting (3.6) into (3.4) gives
@@
where we used the Itô isometry on the second equality. This validates (3.7). The derivatives (3.8) follow directly from definition. A similar argument also appears in [57].
@@
Motivated by Proposition 3.2, we define
@@
Step 1. We compute the derivative of $\mathcal{L}_{0}^{\tau}$ in $\tau$ first. By definition (3.11b),
@@
To estimate the first term, observe that
@@
$=C\left[\beta_{\mu}^{2}\mathcal{W}_{2}\left(\mu^{\tau},\rho^{\mu^{\tau},\alpha^{\tau}}\right)^{2}+\beta_{\mu}^{2}W_{2}\left(\mu^{\tau}_{T},\rho^{\mu^{\tau},\alpha^{\tau}}_{T}\right)^{2}+\beta_{a}^{2}\left\|\nabla_{\alpha}H(t,x,\mu_{t},\alpha^{\tau}(t,x),-\mathcal{G}^{\tau}(t,x))\right\|^{2}_{\mu^{\tau},\alpha^{\tau}}\right],$
where the second inequality follows from (A.27) in Lemma A.13, and the last equality is due to Lemma A.15 and (3.11a). Substituting (C.4) into (C.3) gives
@@
Later, we shall set $\beta_{c}$ sufficiently large relative to $\beta_{a}$ and $\beta_{\mu}$ (cf. (4.11)), so that the positive terms are offset by the decay of the other Lyapunov functions. A similar idea was applied in [58].
Step 2. We analyze $\mathcal{L}^{\tau}_{1}$ next (cf. (C.2)). We treat $\mathcal{L}_{1}$ as a functional of $\mu^{\tau}$, $\alpha^{\tau}$, and $\mathcal{G}^{\tau}$, and define
$\widetilde{\mathcal{L}}_{1}(\mu,\alpha,\mathcal{G})=\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho^{\mu,\alpha}(t,x)\left|\sigma(t,x,\mu_{t})^{\top}\left(\nabla_{x}V^{\mu,\alpha}(t,x)-\mathcal{G}(t,x)\right)\right|^{2}\,\mathrm{d}x\,\mathrm{d}t.$
Then, $\mathcal{L}_{1}^{\tau}=\widetilde{\mathcal{L}}_{1}(\mu^{\tau},\alpha^{\tau},\mathcal{G}^{\tau})$, and the derivative $\partial_{\tau}\mathcal{L}^{\tau}_{1}$ is decomposed into two terms:
@@
where $(c\text{I})$ takes care of the derivative in $\mathcal{G}^{\tau}$ and $(c\text{II})$ deals with the dependence on $(\mu^{\tau},\alpha^{\tau})$. From the flow equation (3.11c), $(c\text{I})$ satisfies
$-(c\text{I})$ $=\beta_{c}\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho^{\mu^{\tau},\alpha^{\tau}}(t,x)\left(\nabla_{x}V^{\mu^{\tau},\alpha^{\tau}}-\mathcal{G}^{\tau}\right)^{\top}D(t,x,\mu^{\tau}_{t})^{2}\left(\nabla_{x}V^{\mu^{\tau},\alpha^{\tau}}-\mathcal{G}^{\tau}\right)\,\mathrm{d}x\,\mathrm{d}t$ (C.7)
$\geq\beta_{c}\sigma_{0}\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho^{\mu^{\tau},\alpha^{\tau}}(t,x)\left(\nabla_{x}V^{\mu^{\tau},\alpha^{\tau}}-\mathcal{G}^{\tau}\right)^{\top}D(t,x,\mu^{\tau}_{t})\left(\nabla_{x}V^{\mu^{\tau},\alpha^{\tau}}-\mathcal{G}^{\tau}\right)\,\mathrm{d}x\,\mathrm{d}t$
$=2\beta_{c}\sigma_{0}\mathcal{L}^{\tau}_{1}.$
We estimate $(c\text{II})$ next. Let $x_{t}=X_{t}^{\mu^{\tau},\alpha^{\tau}}$ be the state dynamics under distribution-control pair $(\mu^{\tau},\alpha^{\tau})$. Denote $\sigma_{t}=\sigma(t,x_{t},\mu_{t}^{\tau})$, $p_{t}=\nabla_{x}V^{\mu^{\tau},\alpha^{\tau}}(t,x_{t})$, and $G_{t}=\mathcal{G}^{\tau}(t,x_{t})$. Then we have
@@
Let $\tau^{\prime}>\tau$, and denote by $x^{\prime}_{t}=X_{t}^{\mu^{\tau^{\prime}},\alpha^{\tau^{\prime}}}$ the state dynamics under $(\mu^{\tau^{\prime}},\alpha^{\tau^{\prime}})$ driven by the same Brownian motion with the same initial condition $x_{0}^{\prime}=x_{0}$. Set $\sigma^{\prime}_{t}=\sigma(t,x^{\prime}_{t},\mu_{t}^{\tau^{\prime}})$, $p^{\prime}_{t}=\nabla_{x}V^{\mu^{\tau^{\prime}},\alpha^{\tau^{\prime}}}(t,x^{\prime}_{t})$, and $G^{\prime}_{t}=\mathcal{G}^{\tau}(t,x^{\prime}_{t})$ (note that it is $\mathcal{G}^{\tau}$ instead of $\mathcal{G}^{\tau^{\prime}}$ here). We have
@@
$=\bigg|\lim_{\tau^{\prime}\to\tau}\dfrac{1}{\tau^{\prime}-\tau}\mathbb{E}\Big[\frac{1}{2}\int_{0}^{T}\left(\left|\sigma_{t}^{\prime\top}(p^{\prime}_{t}-G^{\prime}_{t})\right|^{2}-\left|\sigma_{t}^{\top}(p_{t}-G_{t})\right|^{2}\right)\,\mathrm{d}t\Big]\bigg|$
$\leq\lim_{\tau^{\prime}\to\tau}\dfrac{1}{|\tau^{\prime}-\tau|}\mathbb{E}\Big[\frac{1}{2}\int_{0}^{T}\left|(\sigma_{t}^{\prime\top}p^{\prime}_{t}-\sigma_{t}^{\top}p_{t})-(\sigma_{t}^{\prime\top}G^{\prime}_{t}-\sigma_{t}^{\top}G_{t})\right|\cdot\left|\sigma_{t}^{\prime\top}(p^{\prime}_{t}-G^{\prime}_{t})+\sigma_{t}^{\top}(p_{t}-G_{t})\right|\,\mathrm{d}t\Big]$
@@
$\leq\lim_{\tau^{\prime}\to\tau}\dfrac{1}{|\tau^{\prime}-\tau|}\mathbb{E}\Big[\int_{0}^{T}\dfrac{1}{\sigma_{0}\beta_{c}|\tau^{\prime}-\tau|}\left(\left|\sigma_{t}^{\prime\top}p^{\prime}_{t}-\sigma_{t}^{\top}p_{t}\right|+\left|\sigma_{t}^{\prime\top}G^{\prime}_{t}-\sigma_{t}^{\top}G_{t}\right|\right)^{2}+\dfrac{\sigma_{0}\beta_{c}|\tau^{\prime}-\tau|}{4}\left|\sigma_{t}^{\top}(p_{t}-G_{t})\right|^{2}\,\mathrm{d}t\Big]$
$\leq\lim_{\tau^{\prime}\to\tau}\dfrac{1}{|\tau^{\prime}-\tau|^{2}}\mathbb{E}\Big[\int_{0}^{T}\dfrac{2}{\sigma_{0}\beta_{c}}\left(\left|\sigma_{t}^{\prime\top}p^{\prime}_{t}-\sigma_{t}^{\top}p_{t}\right|^{2}+\left|\sigma_{t}^{\prime\top}G^{\prime}_{t}-\sigma_{t}^{\top}G_{t}\right|^{2}\right)\,\mathrm{d}t\Big]+\dfrac{1}{2}\sigma_{0}\beta_{c}\mathcal{L}_{1}^{\tau}.$
By (A.28) in Lemma A.13, the first term in the last line above satisfies
@@
For the second term, we have
$\quad\mathbb{E}\Big[\int_{0}^{T}\left|\sigma_{t}^{\prime\top}G^{\prime}_{t}-\sigma_{t}^{\top}G_{t}\right|^{2}\,\mathrm{d}t\Big]\leq\mathbb{E}\Big[2\int_{0}^{T}\left(\left|\sigma_{t}^{\prime\top}(G^{\prime}_{t}-G_{t})\right|^{2}+\left|(\sigma_{t}^{\prime}-\sigma_{t})^{\top}G_{t}\right|^{2}\right)\,\mathrm{d}t\Big]$ (C.10)
$\leq\mathbb{E}\Big[2\int_{0}^{T}\left(K^{2}|x^{\prime}_{t}-x_{t}|\right)^{2}+\left(K|x^{\prime}_{t}-x_{t}|K(1+|x_{t}|)\right)^{2}\,\mathrm{d}t\Big]$
[…]
where the second inequality is due to Assumption  4.3, and the last inequality is due to (A.4) in Lemma A.1. Combine (C.9) and (C.10) into (C.8), one has
@@
$=\dfrac{2C}{\beta_{c}}\Big[\beta_{\mu}^{2}\mathcal{W}_{2}\left(\mu^{\tau},\rho^{\mu^{\tau},\alpha^{\tau}}\right)^{2}+\beta_{\mu}^{2}W_{2}\left(\mu^{\tau}_{T},\rho^{\mu^{\tau},\alpha^{\tau}}_{T}\right)^{2}$
$\quad+\beta_{a}^{2}\left\|\nabla_{\alpha}H(t,x,\mu_{t},\alpha^{\tau}(t,x),-\mathcal{G}^{\tau}(t,x))\right\|^{2}_{\mu^{\tau},\alpha^{\tau}}\Big]+\dfrac{1}{2}\sigma_{0}\beta_{c}\mathcal{L}_{1}^{\tau},$
where the last equality is due to lemma A.15. Combining (C.7) and (C.11) into (C.6), we obtain
$\partial_{\tau}\mathcal{L}^{\tau}_{1}\leq-\frac{3}{2}\sigma_{0}\beta_{c}\mathcal{L}_{1}^{\tau}+\dfrac{C}{\beta_{c}}\Big[\beta_{\mu}^{2}\mathcal{W}_{2}\left(\mu^{\tau},\rho^{\mu^{\tau},\alpha^{\tau}}\right)^{2}+\beta_{\mu}^{2}W_{2}\left(\mu^{\tau}_{T},\rho^{\mu^{\tau},\alpha^{\tau}}_{T}\right)^{2}$ (C.12)
@@
where $c_{c}=\frac{3}{2}\sigma_{0}$.
@@
Step 1. We first consider the derivative of $\frac{1}{2}d_{\beta}(\mu^{\tau},\rho^{\mu^{\tau},\alpha^{\tau}})^{2}$ with respect to $\tau$. We decompose the full derivative into $3$ terms, corresponding to the dependence through $\mu^{\tau}$ (left of $d_{\beta}(\mu^{\tau},\rho^{\mu^{\tau},\alpha^{\tau}})$), $\mu^{\tau}$ in $\rho^{\mu^{\tau},\alpha^{\tau}}$, and $\alpha^{\tau}$:
$\quad\dfrac{\mathrm{d}}{\mathrm{d}\tau}\frac{1}{2}d_{\beta}(\mu^{\tau},\rho^{\mu^{\tau},\alpha^{\tau}})^{2}=(\mu\text{I})+(\mu\text{II})+(\mu\text{III})$ (D.1)
$=\dfrac{\mathrm{d}}{\mathrm{d}\tau}\frac{1}{2}d_{\beta}(\mu^{\tau},\nu)^{2}\Big|_{\nu=\rho^{\mu^{\tau},\alpha^{\tau}}}+\dfrac{\mathrm{d}}{\mathrm{d}\tau}\frac{1}{2}d_{\beta}(\mu,\rho^{\mu^{\tau},\alpha})^{2}\Big|_{\mu=\mu^{\tau},\alpha=\alpha^{\tau}}+\dfrac{\mathrm{d}}{\mathrm{d}\tau}\frac{1}{2}d_{\beta}(\mu,\rho^{\mu,\alpha^{\tau}})^{2}\Big|_{\mu=\mu^{\tau}}.$
For $(\mu\text{I})$, by [50, Theorem 5.24], we have for each $t$,
$\quad\dfrac{\mathrm{d}}{\mathrm{d}\tau}\frac{1}{2}W_{2}(\mu^{\tau}_{t},\nu_{t})^{2}\Big|_{\nu_{t}=\rho^{\mu^{\tau},\alpha^{\tau}}_{t}}=-\beta_{\mu}\int_{\mathbb{R}^{d}}|\nabla_{x}\varphi^{\tau}_{t}|^{2}\mathrm{d}\mu^{\tau}_{t}$
@@
Integrating over $t$ with exponential weight gives
$(\mu\text{I})=-\beta_{\mu}\int_{0}^{T}e^{-2\beta t}\,W_{2}(\mu^{\tau}_{t},\rho^{\mu^{\tau},\alpha^{\tau}}_{t})^{2}\,\mathrm{d}t=-\beta_{\mu}d_{\beta}(\mu,\rho^{\mu,\alpha^{\tau}})^{2}.$ (D.2)
Next, we consider $(\mu\text{II})$. By definition,
@@
where the last inequality is due to Lemma A.14. The last term can be simplified through
$\quad\dfrac{\mathrm{d}}{\mathrm{d}\tau}d_{\beta}(\mu^{\tau},\nu)\Big|_{\nu=\mu^{\tau}}=\lim_{\Delta\tau\to 0}\dfrac{1}{\Delta\tau}\Big[\Big(\int_{0}^{T}e^{-2\beta t}\,W_{2}(\mu^{\tau+\Delta\tau},\mu^{\tau})^{2}\,\mathrm{d}t\Big)^{\frac{1}{2}}-0\Big]$ (D.4)
$=\Big[\int_{0}^{T}e^{-2\beta t}\,\Big(\lim_{\Delta\tau\to 0}\dfrac{1}{\Delta\tau}W_{2}(\mu^{\tau+\Delta\tau},\mu^{\tau})\Big)^{2}\,\mathrm{d}t\Big]^{\frac{1}{2}}=\Big[\int_{0}^{T}e^{-2\beta t}\,\Big(\dfrac{\mathrm{d}}{\mathrm{d}\tau}W_{2}(\mu^{\tau},\nu_{t})\Big|_{\nu_{t}=\mu^{\tau}_{t}}\Big)^{2}\,\mathrm{d}t\Big]^{\frac{1}{2}}$
$=\Big[\int_{0}^{T}e^{-2\beta t}\,\left(\beta_{\mu}W_{2}(\mu^{\tau}_{t},\rho^{\mu^{\tau},\alpha^{\tau}}_{t})\right)^{2}\,\mathrm{d}t\Big]^{\frac{1}{2}}=\beta_{\mu}d_{\beta}(\mu^{\tau},\rho^{\mu^{\tau},\alpha^{\tau}}),$
where we used Lemma A.15 in the second last equality. Substitute (D.4) into (D.3), we get
@@
For $(\mu\text{III})$, similar to the estimation for $(\mu\text{II})$, we deduce
@@
$=d_{\beta}(\mu^{\tau},\rho^{\mu^{\tau},\alpha^{\tau}})\left[\lim_{\Delta\tau\to 0^{+}}\dfrac{1}{\Delta\tau^{2}}d_{\beta}(\rho^{\mu^{\tau},\alpha^{\tau+\Delta\tau}},\rho^{\mu^{\tau},\alpha^{\tau}})\right]^{\frac{1}{2}}.$
Next, by (A.5) in Corollary A.2 and (3.11a), (D.6) becomes
@@
Combining (D.2), (D.5), and (D.7) into (D.1), and note that $\kappa\leq\frac{1}{4}$ we obtain
@@
Step 2. Next, we estimate the derivative of $\frac{1}{2}W_{2}(\mu^{\tau}_{T},\rho^{\mu^{\tau},\alpha^{\tau}}_{T})^{2}$. Decompose its derivative in $\tau$ into two terms:
@@
addressing the dependence through $\mu^{\tau}_{T}$, and $\rho^{\mu^{\tau},\alpha^{\tau}}_{T}$, respectively. By [3, Theorem 7.2.2],
$(\mu\text{IV})=-\beta_{\mu}\int\left|\nabla\varphi^{\tau}_{T}\right|^{2}\,\mathrm{d}\mu^{\tau}_{T}=-\beta_{\mu}\int\left|x-T^{\tau}_{T}(x)\right|^{2}\,\mathrm{d}\mu^{\tau}_{T}(x)=-\beta_{\mu}W_{2}(\mu^{\tau}_{T},\rho^{\mu^{\tau},\alpha^{\tau}}_{T})^{2}.$ (D.10)
Similar to the analysis for $(\mu\text{II})$, we have

### v2

@@ Appendix C Proofs for the critic
Proof of Proposition 3.2.
Substituting (3.6) into (3.4) yields
@@
where the second equality follows from the Itô isometry. This validates (3.7) and the derivatives (3.8) follow directly from the definition. Note that a similar argument also appears in [58]. ∎
@@
Proof of Theorem 4.5.
Motivated by Proposition 3.2, define
@@
Step 1. We bound the derivative of $\mathcal{L}_{0}^{\tau}$ in $\tau$. By definition (3.11b),
@@
By (A.27) from Lemma A.13, Lemma A.15 and (3.11a),
@@
$=C\left[\beta_{\mu}^{2}\mathcal{W}_{2}\left(\mu^{\tau},\rho^{\mu^{\tau},\alpha^{\tau}}\right)^{2}+\beta_{\mu}^{2}W_{2}\left(\mu^{\tau}_{T},\rho^{\mu^{\tau},\alpha^{\tau}}_{T}\right)^{2}+\beta_{a}^{2}\left\|\nabla_{\alpha}H(t,x,\mu_{t},\alpha^{\tau}(t,x),-\mathcal{G}^{\tau}(t,x))\right\|^{2}_{\mu^{\tau},\alpha^{\tau}}\right].$
Substituting (C.4) into (C.3) yields
@@
Later, we will set $\beta_{c}$ sufficiently large relative to $\beta_{a}$ and $\beta_{\mu}$ (cf. (4.11)), so that the positive terms are offset by the decay of the other Lyapunov functions. A similar idea was applied in [59].
Step 2. Next, we bound the derivative of $\mathcal{L}^{\tau}_{1}$. We treat $\mathcal{L}_{1}$ as a function of $\mu^{\tau}$, $\alpha^{\tau}$, and $\mathcal{G}^{\tau}$, and define
$\widetilde{\mathcal{L}}_{1}(\mu,\alpha,\mathcal{G}):=\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho^{\mu,\alpha}(t,x)\left|\sigma(t,x,\mu_{t})^{\top}\left(\nabla_{x}V^{\mu,\alpha}(t,x)-\mathcal{G}(t,x)\right)\right|^{2}\,\mathrm{d}x\,\mathrm{d}t,$
so that $\mathcal{L}_{1}^{\tau}=\widetilde{\mathcal{L}}_{1}(\mu^{\tau},\alpha^{\tau},\mathcal{G}^{\tau})$. The derivative $\partial_{\tau}\mathcal{L}^{\tau}_{1}$ is decomposed into two parts:
@@
where $(c\text{I})$ takes care of the $\tau$-dependence through $\mathcal{G}^{\tau}$ and $(c\text{II})$ deals with the $\tau$-dependence through $(\mu^{\tau},\alpha^{\tau})$. From the flow equation (3.11c), $(c\text{I})$ satisfies
$-(c\text{I})$ $=4\beta_{c}\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho^{\mu^{\tau},\alpha^{\tau}}(t,x)\left(\nabla_{x}V^{\mu^{\tau},\alpha^{\tau}}-\mathcal{G}^{\tau}\right)^{\top}D(t,x,\mu^{\tau}_{t})^{2}\left(\nabla_{x}V^{\mu^{\tau},\alpha^{\tau}}-\mathcal{G}^{\tau}\right)\,\mathrm{d}x\,\mathrm{d}t$ (C.7)
$\geq 4\beta_{c}\sigma_{0}\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho^{\mu^{\tau},\alpha^{\tau}}(t,x)\left(\nabla_{x}V^{\mu^{\tau},\alpha^{\tau}}-\mathcal{G}^{\tau}\right)^{\top}D(t,x,\mu^{\tau}_{t})\left(\nabla_{x}V^{\mu^{\tau},\alpha^{\tau}}-\mathcal{G}^{\tau}\right)\,\mathrm{d}x\,\mathrm{d}t$
$=4\beta_{c}\sigma_{0}\mathcal{L}^{\tau}_{1}.$
For part $(c\text{II})$, let $x_{t}:=X_{t}^{\mu^{\tau},\alpha^{\tau}}$ be the state process under $(\mu^{\tau},\alpha^{\tau})$. Denote $\sigma_{t}:=\sigma(t,x_{t},\mu_{t}^{\tau})$, $p_{t}:=\nabla_{x}V^{\mu^{\tau},\alpha^{\tau}}(t,x_{t})$ and $G_{t}:=\mathcal{G}^{\tau}(t,x_{t})$. We have
@@
For $\tau^{\prime}>\tau$, denote by $x^{\prime}_{t}:=X_{t}^{\mu^{\tau^{\prime}},\alpha^{\tau^{\prime}}}$ the state process under $(\mu^{\tau^{\prime}},\alpha^{\tau^{\prime}})$ driven by the same Brownian motion, starting from the same initial condition $x_{0}^{\prime}\overset{\mathrm{a.s.}}{=}x_{0}$. Denote $\sigma^{\prime}_{t}:=\sigma(t,x^{\prime}_{t},\mu_{t}^{\tau^{\prime}})$, $p^{\prime}_{t}:=\nabla_{x}V^{\mu^{\tau^{\prime}},\alpha^{\tau^{\prime}}}(t,x^{\prime}_{t})$ and $G^{\prime}_{t}:=\mathcal{G}^{\tau}(t,x^{\prime}_{t})$. It is worth noting that $G^{\prime}_{t}$ uses $\mathcal{G}^{\tau}$ instead of $\mathcal{G}^{\tau^{\prime}}$. Then,
@@
$=\frac{1}{2}\bigg|\lim_{\tau^{\prime}\to\tau}\dfrac{1}{\tau^{\prime}-\tau}\mathbb{E}\Big[\int_{0}^{T}\left(\left|\sigma_{t}^{\prime\top}(p^{\prime}_{t}-G^{\prime}_{t})\right|^{2}-\left|\sigma_{t}^{\top}(p_{t}-G_{t})\right|^{2}\right)\,\mathrm{d}t\Big]\bigg|$
$\leq\frac{1}{2}\lim_{\tau^{\prime}\to\tau}\dfrac{1}{|\tau^{\prime}-\tau|}\mathbb{E}\Big[\int_{0}^{T}\left|(\sigma_{t}^{\prime\top}p^{\prime}_{t}-\sigma_{t}^{\top}p_{t})-(\sigma_{t}^{\prime\top}G^{\prime}_{t}-\sigma_{t}^{\top}G_{t})\right|\cdot\left|\sigma_{t}^{\prime\top}(p^{\prime}_{t}-G^{\prime}_{t})+\sigma_{t}^{\top}(p_{t}-G_{t})\right|\,\mathrm{d}t\Big]$
@@
$\leq\lim_{\tau^{\prime}\to\tau}\dfrac{1}{|\tau^{\prime}-\tau|}\mathbb{E}\Big[\int_{0}^{T}\dfrac{2}{\sigma_{0}\beta_{c}|\tau^{\prime}-\tau|}\left(\left|\sigma_{t}^{\prime\top}p^{\prime}_{t}-\sigma_{t}^{\top}p_{t}\right|+\left|\sigma_{t}^{\prime\top}G^{\prime}_{t}-\sigma_{t}^{\top}G_{t}\right|\right)^{2}+\dfrac{\sigma_{0}\beta_{c}|\tau^{\prime}-\tau|}{2}\left|\sigma_{t}^{\top}(p_{t}-G_{t})\right|^{2}\,\mathrm{d}t\Big]$
$\leq\lim_{\tau^{\prime}\to\tau}\dfrac{1}{|\tau^{\prime}-\tau|^{2}}\mathbb{E}\Big[\int_{0}^{T}\dfrac{4}{\sigma_{0}\beta_{c}}\left(\left|\sigma_{t}^{\prime\top}p^{\prime}_{t}-\sigma_{t}^{\top}p_{t}\right|^{2}+\left|\sigma_{t}^{\prime\top}G^{\prime}_{t}-\sigma_{t}^{\top}G_{t}\right|^{2}\right)\,\mathrm{d}t\Big]+\sigma_{0}\beta_{c}\mathcal{L}_{1}^{\tau}.$
By (A.28) from Lemma A.13,
@@
By Assumption 4.3 and (A.4) from Lemma A.1,
$\quad\mathbb{E}\Big[\int_{0}^{T}\left|\sigma_{t}^{\prime\top}G^{\prime}_{t}-\sigma_{t}^{\top}G_{t}\right|^{2}\,\mathrm{d}t\Big]\leq 2\mathbb{E}\Big[\int_{0}^{T}\left(\left|\sigma_{t}^{\prime\top}(G^{\prime}_{t}-G_{t})\right|^{2}+\left|(\sigma_{t}^{\prime}-\sigma_{t})^{\top}G_{t}\right|^{2}\right)\,\mathrm{d}t\Big]$ (C.10)
$\leq 2\mathbb{E}\Big[\int_{0}^{T}\left(K^{2}|x^{\prime}_{t}-x_{t}|\right)^{2}+\left(K[|x^{\prime}_{t}-x_{t}|+W_{2}(\mu^{\tau}_{t},\mu^{\tau^{\prime}}_{t})]K(1+|x_{t}|)\right)^{2}\,\mathrm{d}t\Big]$
$\leq C\left(\mathcal{W}_{2}(\mu^{\tau^{\prime}},\mu^{\tau})^{2}+\left\|\alpha^{\tau^{\prime}}-\alpha^{\tau}\right\|_{\mu^{\tau},\alpha^{\tau}}^{2}\right).$
Substituting (C.9) and (C.10) into (C.8) yields
@@
$=\dfrac{C}{\beta_{c}}\Big[\beta_{\mu}^{2}\mathcal{W}_{2}\left(\mu^{\tau},\rho^{\mu^{\tau},\alpha^{\tau}}\right)^{2}+\beta_{\mu}^{2}W_{2}\left(\mu^{\tau}_{T},\rho^{\mu^{\tau},\alpha^{\tau}}_{T}\right)^{2}$
$\quad+\beta_{a}^{2}\left\|\nabla_{\alpha}H(t,x,\mu_{t},\alpha^{\tau}(t,x),-\mathcal{G}^{\tau}(t,x))\right\|^{2}_{\mu^{\tau},\alpha^{\tau}}\Big]+\sigma_{0}\beta_{c}\mathcal{L}_{1}^{\tau},$
where Lemma A.15 is applied. Substituting (C.7) and (C.11) into (C.6) yields
$\partial_{\tau}\mathcal{L}^{\tau}_{1}\leq-3\sigma_{0}\beta_{c}\mathcal{L}_{1}^{\tau}+\dfrac{C}{\beta_{c}}\Big[\beta_{\mu}^{2}\mathcal{W}_{2}\left(\mu^{\tau},\rho^{\mu^{\tau},\alpha^{\tau}}\right)^{2}+\beta_{\mu}^{2}W_{2}\left(\mu^{\tau}_{T},\rho^{\mu^{\tau},\alpha^{\tau}}_{T}\right)^{2}$ (C.12)
@@
which concludes the proof. ∎
@@
Proof of Theorem 4.6.
We bound the derivative $\partial_{\tau}\mathcal{L}^{\tau}_{\mu}$ by taking two steps.
Step 1. We bound the derivative of $\frac{1}{2}d_{\beta}(\mu^{\tau},\rho^{\mu^{\tau},\alpha^{\tau}})^{2}$ with respect to $\tau$, which is further decomposed into $3$ terms that address different sources of $\tau$-dependence: the dependence on $\mu^{\tau}$ through the first argument of $d_{\beta}(\mu^{\tau},\rho^{\mu^{\tau},\alpha^{\tau}})$, the dependence on $\mu^{\tau}$ through the density $\rho^{\mu^{\tau},\alpha^{\tau}}$, and the dependence on $\alpha^{\tau}$ through $\rho^{\mu^{\tau},\alpha^{\tau}}$:
$\quad\frac{1}{2}\dfrac{\mathrm{d}}{\mathrm{d}\tau}d_{\beta}(\mu^{\tau},\rho^{\mu^{\tau},\alpha^{\tau}})^{2}=(\mu\text{I})+(\mu\text{II})+(\mu\text{III})$ (D.1)
$=\frac{1}{2}\dfrac{\mathrm{d}}{\mathrm{d}\tau}d_{\beta}(\mu^{\tau},\nu)^{2}\Big|_{\nu=\rho^{\mu^{\tau},\alpha^{\tau}}}+\frac{1}{2}\dfrac{\mathrm{d}}{\mathrm{d}\tau}d_{\beta}(\mu,\rho^{\mu^{\tau},\alpha})^{2}\Big|_{\mu=\mu^{\tau},\alpha=\alpha^{\tau}}+\frac{1}{2}\dfrac{\mathrm{d}}{\mathrm{d}\tau}d_{\beta}(\mu,\rho^{\mu,\alpha^{\tau}})^{2}\Big|_{\mu=\mu^{\tau}}.$
For $(\mu\text{I})$, by [50, Theorem 5.24], for any $t\in[0,T]$,
$\quad\frac{1}{2}\dfrac{\mathrm{d}}{\mathrm{d}\tau}W_{2}(\mu^{\tau}_{t},\nu_{t})^{2}\Big|_{\nu_{t}=\rho^{\mu^{\tau},\alpha^{\tau}}_{t}}=-\beta_{\mu}\int_{\mathbb{R}^{d}}|\nabla_{x}\varphi^{\tau}_{t}(x)|^{2}\mathrm{d}\mu^{\tau}_{t}(x)$
@@
Multiplying $e^{-2\beta t}$ and integrating both sides with respect to $t$ yield
$(\mu\text{I})=-\beta_{\mu}\int_{0}^{T}e^{-2\beta t}\,W_{2}(\mu^{\tau}_{t},\rho^{\mu^{\tau},\alpha^{\tau}}_{t})^{2}\,\mathrm{d}t=-\beta_{\mu}d_{\beta}(\mu^{\tau},\rho^{\mu^{\tau},\alpha^{\tau}})^{2}.$ (D.2)
Next, we estimate $(\mu\text{II})$. By definition,
@@
where the last inequality follows from Lemma A.14. By Lemma A.15,
$\quad\dfrac{\mathrm{d}}{\mathrm{d}\tau}d_{\beta}(\mu^{\tau},\nu)\Big|_{\nu=\mu^{\tau}}=\lim_{\Delta\tau\to 0}\dfrac{1}{\Delta\tau}\Big[\Big(\int_{0}^{T}e^{-2\beta t}\,W_{2}(\mu^{\tau+\Delta\tau}_{t},\mu^{\tau}_{t})^{2}\,\mathrm{d}t\Big)^{\frac{1}{2}}-0\Big]$ (D.4)
$=\Big[\int_{0}^{T}e^{-2\beta t}\,\Big(\lim_{\Delta\tau\to 0}\dfrac{1}{\Delta\tau}W_{2}(\mu^{\tau+\Delta\tau}_{t},\mu^{\tau}_{t})\Big)^{2}\,\mathrm{d}t\Big]^{\frac{1}{2}}=\Big[\int_{0}^{T}e^{-2\beta t}\,\Big(\dfrac{\mathrm{d}}{\mathrm{d}\tau}W_{2}(\mu^{\tau}_{t},\nu_{t})\Big|_{\nu_{t}=\mu^{\tau}_{t}}\Big)^{2}\,\mathrm{d}t\Big]^{\frac{1}{2}}$
$=\Big[\int_{0}^{T}e^{-2\beta t}\,\left(\beta_{\mu}W_{2}(\mu^{\tau}_{t},\rho^{\mu^{\tau},\alpha^{\tau}}_{t})\right)^{2}\,\mathrm{d}t\Big]^{\frac{1}{2}}=\beta_{\mu}d_{\beta}(\mu^{\tau},\rho^{\mu^{\tau},\alpha^{\tau}}).$
Substituting (D.4) into (D.3) yields
@@
For part $(\mu\text{III})$, we carry out similar estimation to $(\mu\text{II})$:
@@
$=d_{\beta}(\mu^{\tau},\rho^{\mu^{\tau},\alpha^{\tau}})\left[\lim_{\Delta\tau\to 0^{+}}\dfrac{1}{\Delta\tau^{2}}d_{\beta}(\rho^{\mu^{\tau},\alpha^{\tau+\Delta\tau}},\rho^{\mu^{\tau},\alpha^{\tau}})^{2}\right]^{\frac{1}{2}}.$
By (A.5) in Corollary A.2 and (3.11a),
@@
Substituting (D.2), (D.5), and (D.7) into (D.1), and using $\kappa\leq\frac{1}{4}$, we obtain
@@
Step 2. We estimate the $\tau$-derivative of $\frac{1}{2}W_{2}(\mu^{\tau}_{T},\rho^{\mu^{\tau},\alpha^{\tau}}_{T})^{2}$, which is decomposed into two terms:
@@
respectively addressing the $\tau$-dependence through $\mu^{\tau}_{T}$ and $\rho^{\mu^{\tau},\alpha^{\tau}}_{T}$. By [3, Theorem 7.2.2],
$(\mu\text{IV})=-\beta_{\mu}\int_{\mathbb{R}^{d}}\left|\nabla\varphi^{\tau}_{T}(x)\right|^{2}\,\mathrm{d}\mu^{\tau}_{T}(x)=-\beta_{\mu}\int_{\mathbb{R}^{d}}\left|x-T^{\tau}_{T}(x)\right|^{2}\,\mathrm{d}\mu^{\tau}_{T}(x)=-\beta_{\mu}W_{2}(\mu^{\tau}_{T},\rho^{\mu^{\tau},\alpha^{\tau}}_{T})^{2}.$ (D.10)
Similar to the analysis for $(\mu\text{II})$,